# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ 3.2 Sample Characteristics
The official report of the Tsukishima Survey suggests that approximately nine out of ten factory workers in the skilled-age range were classified as skilled workers. Approximately six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^34] Online Appendix A.5 summarizes this point in detail using the complete survey of the primary school students in the 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^35] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector constituted the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is $33$, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^36] Specific occupations are not available for most heads because they are usually listed as “factory workers” in the occupation section. However, sub-categories available for some heads indicate that they were the typical skilled workers in machinery factories such as lathe operators and finishers. Online Appendix A.6 summarizes the work in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen). In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen). [^37] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30–40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima were materially similar. To summarize, this evidence indicates that the THBS sample is useful for inferring the mean tendency of those households.
@@ 3.3 Aggregation
Notes: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure). Figures 4a and 4b show the daily data between January 12 and November 11, 1919. Figures 4c and 4d show the semi-monthly series calculated using the daily data between January 12 and November 11, 1919. The (adjusted) semi-monthly series defines the 12–26th for the first and 27–11th for the second half. Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between January 15 and October 31, 1919. The unadjusted semi-monthly series uses the 15th of each month as a threshold.
Source: Created by the author using the THBS sample.
It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income. To explore a reasonable method of aggregation, I first overviewed the daily time series on the total income and expenditure. Figure 4a shows that the head’s income was paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st). [^38] The head’s income was the dominant method for earnings because the breadwinner households were typical at that time (Panel B in Table 3). Importantly, Figure 4a indicates that the households expensed most of their income on the payday and spent the rest of the payment little by little until the next payday. In other words, the expenditure never clearly increased just before the paydays. This means that the households made their consumption decisions (schedules until the next payday) based on the income they obtained. Note that calendar (semi-)monthly aggregation cannot reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell. [^39] For example, a calendar semi-month cell in January, for example, the cell from 1st to 14th, starts on the 1st but its payday is set for the 14th. Accordingly, this assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell. This is clearly an unrealistic assumption. Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month. Although this shift did not occur frequently, the calendar month cell might contain substantial interference from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal. Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure induces type II errors when testing consumption smoothing.
@@
To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^40] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January). Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st. [^41] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell. My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f). Online Appendix B.8 shows that these measurement errors in the unadjusted dataset strongly attenuate the estimates.
@@ 3.4 Data Description
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I considered five net income categories—savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also considered the sales of miscellaneous goods and labor earnings from the family members (except for the head) to test the contribution of the sales of small assets and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^42] Online Appendix B.11 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^43] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995). As this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^44] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts rarely occurred among the THBS households. Although family size was included in all regressions, to be conservative, it had little effect on the results. [^45] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. I then used the share of children below primary school age range as an alternative time-varying preference shock variable. Including the child share control variable does not influence the results, and the estimates are materially similar to those from my baseline specification (Online Appendix Table B.6). Panel C of Table 3 shows the summary statistics on the family size variable.
@@
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
@@
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
@@
Panel B-1: Savings, insurance, borrowing, and gifts
@@
      Other family member’s earning 3.74 6.11 289 8.21 16.44 124
@@
Notes:
Panel A: This panel shows the summary statistics of the variables used to test the consumption smoothing for the 33- and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Disposable income is the total income minus tax payments.
Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33- and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of miscellaneous goods such as old newspapers and empty bottles.
Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar.
Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed. As those are lagged variables, the initial period is omitted in all the units. Accordingly, a household with two semi-month observations is excluded.
Source: Created by the author from the THBS sample.
@@
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in nominal income during this period. This trend in earnings was, however, offset by steep inflation caused by expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January captures the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The peaks in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These show that the budget data from the THBS households can capture a substantial aspect of the macroeconomic trends.
@@
Online Appendix Figure B.11 illustrates the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. [^46] Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers experienced idiosyncratic income shocks in some cells. Regardless of favorable or adverse shocks, these cells that deviated from the mean tendency help investigate how those households coped with the idiosyncratic shocks. Despite this, it is useful to consider illness as it is the least predictable shock (Gertler and Gruber 2002). Fortunately, the illness among the THBS families is documented systematically and thus, available for analysis. As explained in Section 2.3, neither a comprehensive health insurance system nor compensation for illness from the factories was available. Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates. [^47] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan. The official report mentioned that the workers in the machinery factories in Tsukishima faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p. 14). Although unemployment might be another shock, the THBS heads rarely changed their occupation during the survey period as they were skilled workers who were less likely to move to other jobs. [^48] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey. However, I confirmed that the households with longer and shorter time-series observations had similar family characteristics (Online Appendix B.3). Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk. Moreover, the number of dismissals at the Tsukishima factories would be small. In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory available from the official report of the Tsukishima Survey was only $1.5$% (Department of Health, Ministry of the Interior 1923a, p. 415). Panel D of Table 3 shows the summary statistics on the health shock variables.
@@
The different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). In contrast, this study investigates consumption-smoothing strategies among homogeneous households — skilled factory workers’ households in the machinery manufacturing sector in Tsukishima. To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks, following Schulhofer-Wohl (2011). Specifically, I consider the case that the heads who worked in large-scale or governmental factories might have had different risk preferences because those workers might have had lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are statistically similar. [^49] Online Appendix B.12 summarizes the finer details of the analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and governmental factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery manufacturing sectors had different risk preferences. [^50] The skilled workers in the smithing sector could be business owners (Online Appendix A.6). Similarly, the workers in the non-machinery manufacturing sectors in Tsukishima—a representative machinery production area—could have a different, presumably higher, risk tolerance. This shows that it is plausible to assume that the THBS households shared similar risk preferences, conditional on the household fixed effects.

### v14

@@ 3.2 Sample Characteristics
The official Tsukishima Survey report suggests that approximately nine out of ten factory workers in the skilled-age range are classified as skilled workers. Approximately six of these skilled workers are employed in medium- to large-scale factories, and three in small-scale factories. [^25] Online Appendix A.5 provides detailed evidence using the complete survey of primary school students in Tsukishima. Accordingly, the THBS is designed to examine household budgets of skilled workers. The average age of THBS household heads is $33$, which falls within the representative age range of skilled workers in machinery factories (Kitazawa 1924). [^26] Specific occupations are not available for most heads, as they are generally listed as “factory workers.” However, sub-categories available for some heads indicate typical skilled machinery factory roles such as lathe operators and finishers. Online Appendix A.6 summarizes work in the machinery factories in Tsukishima. The background and rationale for focusing on skilled factory workers in this survey are detailed in Online Appendix B.1. Column 1 of Panel C in Table 2 shows that the average monthly earnings of THBS heads is $59$ yen (median = $56$ yen). Column 2 reports the average monthly earnings of male machinery factory workers in the skilled-age range, calculated using the manufacturing censuses, which is similar at $56$ yen. [^27] This figure is calculated using the average daily wage of male factory workers aged 30–40 and the average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides detailed calculation steps. Average wages for male factory workers follow an inverted-U-shaped distribution with respect to age, peaking in the 30s (Online Appendix Figure B.7). These results indicate that the monthly earnings of THBS household heads closely match those of skilled workers’ households in Tsukishima. Overall, this confirms that the THBS sample is suitable for analyzing the mean tendencies of skilled factory workers’ households in Tsukishima.
@@ 3.3 Aggregation
Notes: Figures 3a3c, and 3e present time-series plots of average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figures 3b3d, and 3f show the corresponding time-series plots of average daily, semi-monthly, and unadjusted semi-monthly net income (income minus expenditure). Figures 3a and 3b display daily data from January 12 to November 11, 1919. Figures 3c and 3d show the semi-monthly series calculated from the daily data over the same period. In the adjusted semi-monthly series, the first half covers the 12–26th and the second half the 27–11th. Figures 3e and 3f present the unadjusted semi-monthly series calculated from daily data between January 15 and October 31, 1919, using the 15th of each month as the threshold. Source: Created by the author using the THBS sample.
Summarizing daily budget information over a longer time span is necessary because households need time to smooth their consumption given their realized income. Figure 3a shows that the head’s income is paid twice a month, typically in the middle (14th–16th) and at the end of the month (30th or 31st). [^28] The head’s income is the primary source of earnings, as most households are typical breadwinner households (Panel B in Table 3). Importantly, Figure 3b indicates that households spend most of their income immediately on payday and gradually use the remainder until the next payday. In other words, expenditure does not increase sharply just before paydays. This suggests that households plan their consumption schedules based on the income they received. Calendar-based (semi-)monthly aggregation cannot capture this behavior because it does not align the payday with the start of the aggregation period. [^29] For example, a calendar semi-month cell in January (1st–14th) starts on the 1st, but the payday falls on the 14th. Aggregating in this way assumes that the consumption within the cell is based on income at the end of the cell, which is unrealistic. Another source of measurement error arises when the payday shifts: occasionally, the payday for the latter half of a month falls on the first day of the next month. While infrequent, this can distort calendar-month aggregation, which may then include three paydays in some months and only one in others. Such mismatches between income and expenditure timing can induce Type II errors when testing for consumption smoothing (Nelson 1994).
@@
To address the timing mismatch, I aggregate daily observations from the 12th to the 26th for the first half of each month and from the 27th to the 11th of the following month for the second half. [^30] This means that the number of days in a semi-month ranges from 13 (February) to 16 (January). This heterogeneity is constant across households, and differences are fully captured by semi-month fixed effects in the regression. This range is based on the observed paydays in the THBS dataset, which fall between the 27th and 31st for the second half of each month. [^31] I do not use a semi-month definition starting on the 2nd (2nd–16th) because it places the payday at the end of the semi-month cell. Online Appendix B.7 shows that although this alternative improves the income series, expenditure cannot correctly follow the income, as expenses measured under this definition are based on income from the previous semi-month. Figure 3c presents the time-series plots of income and expenditure using this adjusted semi-monthly panel, whereas Figure 3e shows the unadjusted (calendar-based) semi-monthly series, using the 15th of each month as the threshold. The unadjusted series clearly provides a rough measure of income and expenditure owing to measurement errors from misaligned paydays. The main improvement in Figure 3c is that expenditure now aligns with income in each month, capturing households’ consumption strategies based on the timing of actual income. As a result, the net income series shows a much smoother trend in Figure 3d, which is not evident in the unadjusted series (Figure 3f). Online Appendix B.8 demonstrates that measurement errors in the unadjusted dataset strongly attenuate the estimates.
@@ 3.4 Data Description
The THBS budget book has two main sections: expenditure and income. Panel A of Table 3 lists the $11$ expenditure subcategories: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These subcategories are used to test the smoothness of the consumptions. To examine the role of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchases, and gifts. Net variables are defined as income minus expenditure in each category; for example, net savings refer to withdrawals minus deposits to savings. I also include the sales of miscellaneous goods and family members’ (excluding the household head’s) labor earnings to assess the contribution of small asset sales and intrahousehold labor supply adjustments. Panel B of Table 3 summarizes these variables. [^32] Online Appendix B.11 presents the results of panel unit root tests. The null hypothesis of unit roots is rejected at conventional levels for all these variables.
Changes in family size may affect household preferences (Jappelli and Pistaferri 2017). [^33] In village economies, landholding and assets such as livestock and grain may also correlate with households’ risk-coping strategies (e.g., Udry 1994, 1995). As this study focuses on urban factory-worker households, the roles of these assets are not relevant. Fortunately, the archives include sheets documenting changes in the composition of THBS households (Kazoku idō hyō). [^34] OISR, Archives of the Tsukishima Survey (THBS, unreleased). I find that household size remains stable over time. During the sample period, only one household experienced a change in family size due to the birth of a son. This indicates that preference shifts rarely occurred among THBS households. Although family size is included in all regressions, it has minimal effect on the results. [^35] The family size variable is omitted in some regressions owing to multicollinearity in fixed-effects models, given its small within-household variation. As an alternative time-varying measure of preference shocks, I use the share of children below primary school age. Including this child-share variable does not affect the results; estimates are materially similar to the baseline specification (Online Appendix Table B.6). Panel C of Table 3 presents the summary statistics for family size.
@@
Semi-monthly panel dataset Monthly panel dataset
@@
Semi-monthly panel dataset Monthly panel dataset
@@
Panel B-1: Net savings, insurance, borrowing, and gifts (income - expenditure)
@@
    Other family members’ earnings 3.74 6.11 289 8.21 16.44 124
@@
Notes: Panel A: This panel presents summary statistics for the variables used to test consumption smoothing for the 33- and 26-panel units in the adjusted semi-monthly and adjusted monthly panel datasets, respectively. Disposable income is defined as total income minus tax payments. Panel B: This panel presents summary statistics for the variables used to examine risk-coping strategies for the 33- and 26-panel units in the adjusted semi-monthly and adjusted monthly panel datasets, respectively. Each net income variable is defined as income minus expenditure. “Net savings”: withdrawals minus deposits to savings. “Net insurance”: insurance payments received minus expenses paid for insurance. “Net borrowing”: the amount borrowed minus debt repayments. This category includes net income from both pawnshops and other lending institutions. “Net credit purchase”: credit purchases minus credit repayments. “Net gifts”: total gifts received, including both pecuniary and non-pecuniary, minus any payments for gifts. “Other family members’ earnings”: total income earned by all family members except the household head. “Sales of miscellaneous assets”: include the sale of goods such as old newspapers and empty bottles. Panel C: This panel shows summary statistics for the family size variable for the 33-panel units in the adjusted semi-monthly panel dataset. Statistics for the adjusted monthly panel dataset are not reported because they are materially similar. Panel D: This panel shows summary statistics for the number of days family members were ill among the 33-panel units in the adjusted semi-monthly panel dataset. Since these are lagged variables, the initial period is omitted for all units. Accordingly, households with only two semi-month observations are excluded. Source: Created by the author from the THBS sample.
@@
Figure 3c shows a slight upward trend in average income, consistent with the overall macroeconomic increase in nominal income during this period. However, this growth in earnings is offset by steep inflation following postwar expansions (Section 2.1). In Figure 3d, the deviation of expenditure from income in January reflects spending on goods for New Year events. The gaps from June to the first half of July likely correspond to bonus payments. Peaks in March may indicate preparation for the new fiscal and academic year starting in April, suggesting a higher share of temporary income from withdrawals and loans. These patterns demonstrate that the THBS household budget data capture key aspects of macroeconomic trends.
@@
Online Appendix Figure B.11 illustrates the residuals from the regression of the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. [^36] Let $n_{t}$ be the number of households in semi-month $t$ in 1919. For household $i$ in semi-month $t$, the residual is defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. Although the residuals satisfy the zero-average property, they occasionally take large positive or negative values, indicating that workers experienced idiosyncratic income shocks in some cells. These deviations from the mean help examine household responses to idiosyncratic shocks. Fortunately, the THBS archives (Kazoku idō hyō) document the exact timing and duration of family members’ illnesses. To understand the mechanisms behind consumption smoothing, I treat illness as the least predictable adverse shock (Gertler and Gruber 2002). While unemployment could also act as a shock, THBS household heads rarely changed occupations during the survey period, as they were skilled workers unlikely to move to other jobs. [^37] Household heads who changed jobs may have been omitted from the survey. However, households with longer and shorter time-series observations exhibit similar family characteristics (Online Appendix B.3), so this omission does not create selection bias regarding vulnerability to idiosyncratic risk. Moreover, dismissals at the Tsukishima factories were rare. For example, the average monthly dismissal rate from June 1919 to May 1920 at the one factory with available data was 1.51.5% (Department of Health, Ministry of the Interior 1923a, p. 415). Panel D of Table 3 reports summary statistics for the health shock variables.
@@
Different attitudes toward risk across households with varying occupations and in diverse regions can bias standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). However, this study focuses on consumption-smoothing strategies among homogeneous households—skilled factory workers in the machinery manufacturing sector in Tsukishima. To assess the homogeneity of risk preferences among THBS households, I test whether the earnings of household heads in specific factories or sectors responded differently to aggregate shocks, following Schulhofer-Wohl (2011). The results show that the sensitivities to aggregate shocks were statistically similar. [^38] The heads’ earnings are regressed on aggregate consumption, its interaction with an indicator for large-scale or governmental factories (or for smithing and non-machinery factories), the family size variable, and household fixed effects. The results remain unchanged when a few semi-months with smaller cross-sectional samples are trimmed. THBS households shared similar risk preferences, conditional on household fixed effects. Online Appendix B.12 provides additional details of the analysis.